OpenStudy (anonymous):

Simplify (6x^2 - 3 - 5x^3) - (4x^3 + 2x^2 - 8)

OpenStudy (anonymous):

First thing's first Let's simplify step-by-step. 6x2−3−5x3−(4x3+2x2−8) Distribute the Negative Sign: =6x2−3−5x3+−1(4x3+2x2−8) =6x2+−3+−5x3+−1(4x3)+−1(2x2)+(−1)(−8) =6x2+−3+−5x3+−4x3+−2x2+8 Combine Like Terms: =6x2+−3+−5x3+−4x3+−2x2+8 =(−5x3+−4x3)+(6x2+−2x2)+(−3+8) =−9x3+4x2+5 Answer: =−9x3+4x2+5

OpenStudy (anonymous):

thank you